### v2

# Liu and Zhang

[…]
Liu and Zhang
[…]
PCL for RMAB with General Observation Models
[…]
PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models
[…]
In this paper, we consider a general observation model for restless multi-armed bandit problems. The operation of the player needs to be based on certain feedback mechanism that is error-prone due to resource constraints or environmental or intrinsic noises. By establishing a general probabilistic model for dynamics of feedback/observation, we formulate the problem as a restless bandit with a countable belief state space starting from an arbitrary initial belief (a priori information). We apply the achievable region method with partial conservation law (PCL) to the infinite-state problem and analyze its indexability and priority index (Whittle index). Finally, we propose an approximation process to transform the problem into which the AG algorithm of Niño-Mora and Bertsimas for finite-state problems can be applied to. Numerical experiments show that our algorithm has an excellent performance.
[…]
1 Introduction
Multi-armed bandit (MAB) is a classic operations research problem that involves a learner making choices among actions with uncertain/random rewards in order to maximize the expected return in long-run. The MAB problem is often associated with the important exploration-exploitation tradeoff and was initially proposed by Robbins (1952). Starting from the classic stochastic scheduling problem, the ongoing development of various MAB models makes it applicable to a wide range of practical fields, including clinical trials, recommendation systems, cognitive communications, and financial investments (Gittins 1979, Berry and Fristedt 1985, Press 2009, Farias and Madan 2011, Hoffman et al. 2011, Shen et al. 2015).
In the classic Bayesian MAB problem, the bandit machine has a total of $N$ arms and one player pulling an arm in each time slot. After a player selected one of these arms and activated it, a random reward will be accrued depending on the activated arm and its current state. In this process, all states are visible, and only the state of the arm being activated changes according to a Markov chain. The goal of the MAB problem is to search a policy that maximizes the long-term cumulative reward. Gittins (1979) proved that the classic MAB problem can be solved optimally by an index policy, referred to as Gittins index, and the player only needs to activate the arm with the largest index at each moment. The index policy significantly reduces the complexity of the problem from being exponential with the number of arms to being linear. After that, Whittle (1988) extended the MAB problem to the restless MAB (RMAB) model, in which $K$ arms can be activated at each moment, and the arms not chosen may also undergo state transitions over time. However, for the general RMAB problem with a finite state space, Papadimitriou and Tsitsiklis (1994) showed that the computational complexity of finding the optimal policy is already PSPACE-hard. Fortunately, Whittle (1988) generalized Gittins index to Whittle index that provides a possible solution to RMAB by considering Lagrangian relaxation and duality. Accordingly, Whittle index is optimal under a relaxed constraint that the time-average number of activated arms is $K$. For the original problem, the Whittle index policy has been proven to be asymptotically optimal per-arm wise as the number of arms goes to infinity under certain conditions (Weber and Weiss 1990, 1991). Nonetheless, not every RMAB is indexable, i.e., Whittle index may not exist. And the establishment of indexability is itself a difficult problem. Even if indexability is proven for a particular RMAB, analytical solutions of the Whittle index function can still be hard to obtain. For some important categories of RMAB models, Whittle indexability and its strong performance have been demonstrated in the literature, such as the dual-speed bandit problem and partially observable RMAB (Glazebrook et al. 2002, Ahmad et al. 2009, Liu and Zhao 2010, Gittins et al. 2011, Liu 2021, Liu et al. 2024).
In this paper, we extend the partially observable RMAB considered in Liu et al. (2010, 2024) to general observation models. The previous work only considers special classes of observation errors and noises within the framework of partially observation Markov decision processes (POMDP) (Zhao and Sadler 2007, Liu et al. 2010, 2024). But general algorithms for POMDP often suffer from curse of dimensionality and becomes cumbersome when the value functions for dynamic programming are too complex (Smallwood and Sondik 1973, Sondik 1978). In order to deal with general observation models, we need an alternative methodology to design efficient algorithms for systems with large sets of parameters. In the interdisciplinary field of operations research and stochastic optimization, an analytical method called “Achievable Region” emerged in the 1990s. Interestingly, this method transforms a time-series optimization problem into linear programming (LP) problems that can be solved efficiently (Coffman and Mitrani 1980, Federgruen and Groenevelt 1988a, b, Gelenbe and Mitrani 2010). The challenge of using this method is mainly on proving the feasibility of such transformation (Bertsimas 1995). Referred to as the general conservation law (GCL), Bertsimas and Niño-Mora (1996) proposed the required structure of performance measures for the achievable region to be a polyhedron called extended polymatroid. Based on GCL, they extended Klimov’s algorithm to several stochastic scheduling problems, including the classic MAB problem (Klimov 1975, Thomas 1991). By relaxing the GCL restrictions, Niño-Mora (2001) proposed the partial conservation law (PCL) and specified conditions for an RMAB to satisfy PCL and subsequently lead to the numerical calculation of Whittle index. Later on, Niño-Mora (2002) offered an economic explanation of PCL-indexability and its relation to Whittle indexability where Whittle index can be interpreted as the optimal marginal cost rate. A more comprehensive summary of this work can be found in Niño-Mora (2007). However, all such work considers finite state spaces except that Frostig and Weiss (1999) extended the GCL framework for classic MAB to countable state spaces. In this paper, we extend the PCL framework to the RMAB with a general parametrization of observation model and an infinite state space. The main challenge of this extension is due to the high-dimensional probability state space (belief states) with complex dynamics for transitions over time. By establishing the weak duality between LP formulations, we build the PCL framework for analyzing our RMAB and subsequently design an efficient algorithm for calculating Whittle index when PCL-indexability is satisfied. Finally, we demonstrate the superior performance of our algorithm by numerical simulations.
2 Main Results
2.1 Model Formulation
In this section, we begin with the formulation of the RMAB problem with general observation models. Assume the system has $N$ arms and the state space of the $n$th arm is $\mathcal{M}_{n}$. Therefore the entire system state space is $\mathcal{M}=\prod_{n=1}^{N}\mathcal{M}_{n}$. At each moment, the actual state of each arm undergos a state transition based on its own Markov probability transition matrix, and we will select $K$ arms to activate. For those activated arms, we can observe their states (with errors) and accrue some reward dependent on the observed states and the actual states. For those arms that are not activated, we can neither observe their states nor obtain reward from them, and their states still transit over time. Let $S_{n}(t),O_{n}(t)\in\mathbb{Z}^{+}$ be respectively the true and observed states of arm $n$ in slot $t$ and $A(t)$ the set of activated arms in slot $t$. First consider a single-arm process (i.e. $N=1$ and $\mathcal{M}=\mathcal{M}_{n}$) and drop the subscript $n$ (sometimes the time index $t$ is also dropped if referring to the same time index) for convenience. Suppose the state transition matrix, error matrix and reward matrix are $P=\{p_{ij}\}_{\mathcal{M}\times\mathcal{M}}$, $E=\{\varepsilon_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ and $R=\{r_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ respectively. For error matrix, $\varepsilon_{ij}$ represents the probability that the observed state is $j$ when the true state is $i$, that is, $\varepsilon_{ij}=P(O=j|S=i)$, while $r_{ij}$ represents the reward obtained when the true state is $i$ and the observed state is $j$ (different observations may lead to different sub-actions and thus different rewards following). Denoted by ${\Omega}_{a}=\{(\omega_{1},\cdots,\omega_{M})|\sum_{i=1}^{M}\omega_{i}=1,\omega_{i}\geq 0,1\leq i\leq M:=|\mathcal{M}|\}$ the belief space, for each $\boldsymbol{\omega}(t)=(\omega_{1}(t),\cdots,\omega_{M}(t))\in{\Omega_{a}}$, $\omega_{i}(t)$ represents the conditional probability that the true state is $i$ in slot $t$ (based on past observations). If the current belief state is $(\omega_{1},\cdots,\omega_{M})$, the current expected reward is $\sum_{i=1}^{M}\omega_{i}(\sum_{j=1}^{M}\varepsilon_{ij}r_{ij})$. In addition, we may sometimes receive additional feedback related to the true state and the immediate reward obtained that helps us better estimate the actual arm state (e.g., ACK/NAK in a communication channel to indicate whether the data was successfully transmitted (Liu et al. 2024)). Assume that there are $L$ feedback states (positive integers) in total that encompass all possible observations, and denote the feedback state at time $t$ by $F(t)~{}(L\leq M^{2})$. Define $\rho_{ij}=P(F(t)=j|S(t)=i)$. By Bayes rule, we have
$P(S(t+1)=j|F(t)=i)=\frac{P(S(t+1)=j,F(t)=i)}{P(F(t)=i)}=\frac{\sum_{k=1}^{M}p_{kj}\rho_{ki}\omega_{k}}{\sum_{k=1}^{M}\rho_{ki}\omega_{k}}.$ (1)
Thus the rule of belief update is
$\boldsymbol{\omega}(t+1)=\begin{cases}(\frac{\sum_{i=1}^{M}p_{i1}\rho_{ij}\omega_{i}}{\sum_{i=1}^{M}\rho_{ij}\omega_{i}},\cdots,\frac{\sum_{i=1}^{M}p_{iM}\rho_{ij}\omega_{i}}{\sum_{i=1}^{M}\rho_{ij}\omega_{i}}),&\text{if active and }F(t)=j\\ \boldsymbol{\omega}(t)P,&\text{if passive}\end{cases}.$ (2)
In a simpler scenario where there is no additional feedback but only the observed states, the update rule becomes
$\boldsymbol{\omega}(t+1)=\begin{cases}(\frac{\sum_{i=1}^{M}p_{i1}\varepsilon_{ij}\omega_{i}}{\sum_{i=1}^{M}\varepsilon_{ij}\omega_{i}},\cdots,\frac{\sum_{i=1}^{M}p_{iM}\varepsilon_{ij}\omega_{i}}{\sum_{i=1}^{M}\varepsilon_{ij}\omega_{i}}),&\text{if active and }O(t)=j\\ \boldsymbol{\omega}(t)P,&\text{if passive}\end{cases}.$ (3)
In the extreme case when we cannot observe any state after activating an arm, we simply treat the obtained reward (might be 00) as the feedback. By normalizing the immediate reward $r(t)$ to take only positive integer values, we have
$\rho_{ij}=$ $P(r(t)=j|S(t)=i)$
$=$ $\frac{P(r(t)=j,S(t)=i)}{P(S(t)=i)}$
$=$ $\frac{\sum_{k=1}^{M}P(r(t)=j,O(t)=k,S(t)=i)}{\omega_{i}}$
$=$ $\frac{\sum_{k=1}^{M}P(r(t)=j|O(t)=k,S(t)=i)P(O(t)=k|S(t)=i)\omega_{i}}{\omega_{i}}$
$=$ $\sum_{k=1}^{M}\mathbbm{1}(r_{ik}=j)\varepsilon_{ik}$
or equivalently $\rho_{ir}=\sum_{j=1}^{M}\mathbbm{1}(r_{ij}=r)\varepsilon_{ij}$. Thus we have
$\sum_{i=1}^{M}p_{i1}\rho_{ir}\omega_{i}={}$ $\sum_{i=1}^{M}p_{i1}(\sum_{j=1}^{M}\mathbbm{1}(r_{ij}=r)\varepsilon_{ij}\omega_{i})$
[…]
Similarly,
$\sum_{i=1}^{M}\rho_{ir}\omega_{i}=\sum_{r_{ij}=r}\varepsilon_{ij}\omega_{i}.$ (5)
Combining (2.1) and (5), the belief state update rule is
$\boldsymbol{\omega}(t+1)=\begin{cases}(\frac{\sum_{r_{ij}=r}p_{i1}\varepsilon_{ij}\omega_{i}}{\sum_{r_{ij}=r}\varepsilon_{ij}\omega_{i}},\cdots,\frac{\sum_{r_{ij}=r}p_{iM}\varepsilon_{ij}\omega_{i}}{\sum_{r_{ij}=r}\varepsilon_{ij}\omega_{i}}),&\text{if active and }r(t)=r\\ \boldsymbol{\omega}(t)P,&\text{if passive}\end{cases}.$ (6)
A more practical situation is that we can jointly obtain information from the observed state and the obtained reward. In this case, we can similarly give the belief update rule:
$\boldsymbol{\omega}(t+1)=\begin{cases}(\frac{\sum_{i=1}^{M}p_{i1}\varepsilon_{ij}\omega_{i}\mathbbm{1}(r_{ij}=r)}{\sum_{i=1}^{M}\varepsilon_{ij}\omega_{i}\mathbbm{1}(r_{ij}=r)},\cdots,\frac{\sum_{i=1}^{M}p_{iM}\varepsilon_{ij}\omega_{i}\mathbbm{1}(r_{ij}=r)}{\sum_{i=1}^{M}\varepsilon_{ij}\omega_{i}\mathbbm{1}(r_{ij}=r)}),&\text{if active and }O(t)=j,\\ &r(t)=r~{}(\exists i~{}s.t.~{}r_{ij}=r)\\ \boldsymbol{\omega}(t)P,&\text{if passive}\end{cases}.$ (7)
For the restless multi-armed bandit model, the goal is to find a policy $\pi$ which maps the belief states of all arms into an active set $A(t)$ in slot $t$ that maximize the long-term expected discounted reward. In other words, if we denote the reward obtained by the $n$-th arm in slot $t$ by $r_{n}(t)$, then our objective is
$\max_{\pi}\mathbb{E}_{\pi}$ $\left[\sum_{t=1}^{+\infty}\beta^{t-1}\sum_{n=1}^{N}r_{n}(t)\bigg{|}\boldsymbol{\omega}_{1}(1),\cdots,\boldsymbol{\omega}_{n}(1)\right],$ (8)
$s.t.~{}|A(t)|=K,t\geq 1,$ (9)
where $10\leq\beta<$ 1 is the discount factor and $\boldsymbol{\omega}_{n}(t)$ is the belief state (vector) of arm $n$ in slot $t$. In this problem, the diversity of states, choices, errors makes the problem highly complex. In RMAB problems, searching for an easily computable priority index policy is the mainstream. The core idea is to assign an index (a real number) to the current state of each arm, and then activate those arms with top $K$ large indexes. The goal of this paper is to theoretically characterize the conditions when such an index policy exists and provide a detailed algorithm for efficiently computing the index function (if it exists).
2.2 Whittle Index
Whittle (1988) relaxed the constraint on the exact number of arms activated in each slot, requiring only that the expected number of arms activated per slot on average (in the sense of discounted time) is $K$, i.e.
$\mathbb{E}_{\pi}\left[\sum_{t=1}^{+\infty}\beta^{t-1}\sum_{n=1}^{N}\mathbbm{1}(n\in A(t))\bigg{|}\boldsymbol{\Omega}(1)\right]=\frac{K}{1-\beta}$ (10)
or
$\mathbb{E}_{\pi}\left[\sum_{t=1}^{+\infty}\beta^{t-1}\sum_{n=1}^{N}\mathbbm{1}(n\notin A(t))\bigg{|}\boldsymbol{\Omega}(1)\right]=\frac{N-K}{1-\beta},$ (11)
where $\boldsymbol{\Omega}(t)=(\boldsymbol{\omega}_{1}(t),\cdots,\boldsymbol{\omega}_{N}(t))$ with $\boldsymbol{\omega}_{n}(t)$ as the belief state (vector) for arm $n$ at time $t$. Thus the Lagrange optimization problem can be written as
$\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=1}^{+\infty}\beta^{t-1}\sum_{n=1}^{N}(\mathbbm{1}(n\in A(t))r_{n}(t)+\lambda\mathbbm{1}(n\notin A(t)))\bigg{|}\boldsymbol{\Omega}(1)\right].$ (12)
The above problem can be decomposed into $N$ independent subproblems, that is, for any $1\leq n\leq N$,
$\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=1}^{+\infty}\beta^{t-1}(\mathbbm{1}(n\in A(t))r_{n}(t)+\lambda\mathbbm{1}(n\notin A(t)))\bigg{|}\boldsymbol{\omega}_{n}(1)\right].$ (13)
The explanation of the optimization problem above is that for each arm, when it is not selected (made passive), we will receive a subsidy $\lambda$. Since the problems above are independent, we just need to consider the single-arm case. For notation simplicity, we drop the arm index $n$ from now on. For a given arm, the optimal policy for the relaxed optimization problem divides the arm state space into two subset (here the arm state space is the arm belief state space): active set $\mathcal{A}(\lambda)$ and passive set $\mathcal{P}(\lambda)$. Specifically, $\mathcal{P}(\lambda)$ contains all belief states in which the optimal choice is passive when the subsidy is $\lambda$. In particular, for a certain state $s$, if both active and passive actiosn are optimal, we include it in $\mathcal{P}(\lambda)$, and $\mathcal{A}(\lambda)$ is just the complement of $\mathcal{P}(\lambda)$ in the entire state space. Under the concept of passive set, Whittle indexability can be stated as follows:
Definition 2.1 (Whittle Indexability)
An arm is indexable if the passive set $\mathcal{P}(\lambda)$ increases from $\emptyset$ to the whole state space as the subsidy $\lambda$ increases from $\infty$- ∞ to +$\infty$+ ∞. The RMAB problem is indexable if every decoupled problem is indexable.
Indexability states that, once an arm is made passive with subsidy $\lambda$, it should also be passive with any $\lambda^{\prime}$ larger than $\lambda$. If the problem satisfies indexability, its Whittle index is defined as follows:
Definition 2.2 (Whittle Index)
If an arm is indexable, the Whittle index $W(s)$ of a state $s$ is the infimum subsidy $\lambda$ that keeps $s$ in the passive set $\mathcal{P}(\lambda)$. That is,
$W(s):=\inf\{\lambda:s\in\mathcal{P}(\lambda)\}.$
By continuity, $W(s)$ is the infimum subsidy that makes it equivalent to be active or passive at state $s$. Under the definitions of indexability and Whittle index, the Whittle index policy for the original multi-armed bandit problem is simply to activate the $K$ arms whose states offer the largest Whittle indices. In fact, for the classic MAB problem where passive arms do not change states, the Whittle indexability is always satisfied and the Whittle index is reduced to the Gittins index.
2.3 Belief State Space
Different from the perfect observation case, the update of the belief state is nonlinear for the general observation model. This yields much difficulty for us to use value functions and dynamic programming methods to analyze the problem. Given an initial belief state, it appears that the size of the belief state space will grow exponentially over time as all possible realizations of observations/rewards/feedback are incessantly traversed and updated. Fortunately, a large number of numerical experiments have shown that in the sense of Euclidean norm approximation, the state space exhibits convergence after a process of iterative calculations (see Sec. 3 for details). Note that the initial belief state is $\boldsymbol{\omega}$, $\{\mathcal{B}_{l}\}_{0\leq l\leq L}$ is a list of operators defined by feedback states and state update rules, that is, for $1\leq l\leq L$, $\mathcal{B}_{l}\boldsymbol{\omega}$ is the updated state caused by the $l$-th feedback state under the active action, and $\mathcal{B}_{0}$ is the state update operator under the passive action. In this paper, $\mathcal{B}_{0}\boldsymbol{\omega}=\boldsymbol{\omega}P$. Define the $T$-step state space recursively as follows:
Definition 2.3 ($T$-step state space)
Define
$\Omega(T|\boldsymbol{\omega})=\begin{cases}\{\mathcal{B}_{l_{1}}\cdots\mathcal{B}_{l_{T}}\boldsymbol{\omega}:0\leq l_{1},\cdots,l_{T}\leq L\}\cup\Omega(T-1|\boldsymbol{\omega}),&T\geq 2\\ \{\mathcal{B}_{l}\boldsymbol{\omega}:0\leq l\leq L\}\cup\{\boldsymbol{\omega}\},&T=1\end{cases}.$
We call $\Omega(T|\boldsymbol{\omega})$ the $T$-step state space under the initial belief state $\boldsymbol{\omega}$.
Under the definition of the $T$-step state space, $\Omega(+\infty|\boldsymbol{\omega_{0}})$ is the belief space obtained by traversing all possible state update rules starting from the initial state $\boldsymbol{\omega_{0}}$. Clearly, $\Omega(+\infty|\boldsymbol{\omega_{0}})$ is countable and we denote it by $\Omega(\boldsymbol{\omega_{0}})$. Under a policy $\pi$, let $p_{ij}(\pi):=P(\boldsymbol{\omega}(t+1)=\boldsymbol{\omega_{j}}|\boldsymbol{\omega}(t)=\boldsymbol{\omega_{i}})$ be the conditional probability of transition from belief state $\boldsymbol{\omega_{i}}$ to $\boldsymbol{\omega_{j}}$. Due to different actions of activation and passivity affecting the state update, decomposition of $p_{ij}(\pi)$ should be carried out. Let $a(t)$ be the indicator function of whether the arm is activated in slot $t$, then we define the transition probability of the belief state as follows:
$p_{ij}^{0}=P(\boldsymbol{\omega}(t+1)=\boldsymbol{\omega}_{j}|\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i},a(t)=0),$ (14)
$p_{ij}^{1}=P(\boldsymbol{\omega}(t+1)=\boldsymbol{\omega}_{j}|\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i},a(t)=1).$ (15)
By (14)-(15) and the total probability theorem, we have
$p_{ij}(\pi)$ $=\sum\limits_{a=0}^{1}P(a(t)=a|\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i})p_{ij}^{a}.$ (16)
The elements in the probability transition matrix under passive and active actions have the following more specific expressions:
$p_{ij}^{0}=\begin{cases}1,&\text{if $\boldsymbol{\omega}_{j}=\boldsymbol{\omega}_{i}P$}\\ 0,&\text{otherwise}\end{cases}$ (17)
and
$p_{ij}^{1}$ $=P(\boldsymbol{\omega}(t+1)=\boldsymbol{\omega}_{j}|\text{active and }\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i})$ (18)
=$=\sum\limits_{\{l:\mathcal{B}_{l}\boldsymbol{\omega}_{i}=\boldsymbol{\omega}_{j},1\leq l\leq L\}}P(F(t)=l|\text{active and }\boldsymbol{\omega}(t)=\boldsymbol{\omega}_{i}).$
In the following sections, these two probability transition matrices will play an important role in the calculation of Whittle indices.
2.4 Two-Arm Problem and Achievable Region
With the rapid development at the junction of operations research, stochastic optimization and reinforcement learning, quite a few effective methods for finite-state problems have been found, among which achievable region with conservation laws is a great example. Bertsimas and Niño-Mora (1996) and Niño-Mora (2001) adopted the analytical framework of generalized conservation laws and partial conservation laws for the classic MAB and RMAB problems with finite state spaces, respectively, and provided efficient index algorithms for the corresponding problems. In the rest subsections, we will apply the PCL framework to the analysis of our RMAB problem with an infinite state space, and theoretically build the foundation for the construction of an efficient index policy.
Consider the single-arm process with $M$ states discussed in the previous section. For an initial belief state $\boldsymbol{\omega_{0}}$, let $\Omega(\boldsymbol{\omega_{0}})$ be the countable belief state space generated by iteratively updating $\boldsymbol{\omega_{0}}$ through state transitions. During the process of making the arm active or passive, since all belief states fall within $\Omega(\boldsymbol{\omega_{0}})$, the entire time period can be completely partitioned by the time segments in which each state is located. In this scenario, define the performance indicator variables for each belief state as follows:
$I_{\boldsymbol{\omega}}^{1}(t)=\begin{cases}1,&\text{if the arm is active in slot $t$ and its belief state is $\boldsymbol{\omega}$}\\ 0,&\text{otherwise}\end{cases}$
and
$I_{\boldsymbol{\omega}}^{0}(t)=\begin{cases}1,&\text{if the arm is passive in slot $t$ and its belief state is $\boldsymbol{\omega}$}\\ 0,&\text{otherwise}\end{cases}.$
Furthermore, define the performance measures of belief state $\boldsymbol{\omega}$ as follows:
$x_{\omega}^{a}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{+\infty}\beta^{t}I_{\boldsymbol{\omega}}^{a}(t)\right],\quad a=0,1$ (19)
$x_{\omega}^{a}(\pi,\boldsymbol{\omega}_{i})=\mathbb{E}_{\pi}\left[\sum_{t=0}^{+\infty}\beta^{t}I_{\boldsymbol{\omega}}^{a}(t)\bigg{|}\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad a=0,1$ (20)
where $\boldsymbol{\omega}(t)$ is the belief state of the arm in slot $t$. In Whittle’s relaxation, the Lagrangian multiplier $\lambda$ can be regarded as a subsidy when the arm is made passive. Thus the optimization objective of the partially observable RMAB problem with subsidy can be written as follows:
$\max_{\pi}$ $\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}+\lambda\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega}}^{0}$ (21)
subject to $x_{\boldsymbol{\omega}_{i}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0}=e_{\boldsymbol{\omega}_{i}}+\beta(\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega}_{j}}^{1}+\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega}_{j}}^{0}),\quad\forall\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$
$x_{\boldsymbol{\omega}_{i}}^{1},x_{\boldsymbol{\omega}_{i}}^{0}\geq 0,\quad\forall\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$
where $R_{\boldsymbol{\omega}}=\sum\limits_{i=1}^{M}\omega_{i}R_{i},\boldsymbol{\omega}=(\omega_{1},\cdots,\omega_{M})$, $p_{ij}^{a}$ are given by (17)-(18), $e_{\boldsymbol{\omega}_{i}}$ is the indicator whether the initial belief state is $\boldsymbol{\omega}_{i}$, $r_{i}$ is the expected immediate reward when $S(t)=i$ with the active action. Obviously, $R_{\boldsymbol{\omega}}$ is bounded above and we can assume $|R_{\boldsymbol{\omega}}|\leq C,\forall\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})$. For the equality constraints of this optimization problem, it can be understood that the left side of the equation is the direct performance measure for state $\boldsymbol{\omega}_{i}$, while the right side of the equation is another way to represent it. That is, each occurrence of state $\boldsymbol{\omega}_{i}$ in the system is transitioned by another state $\boldsymbol{\omega}_{j}$. Therefore, the performance measure for state $\boldsymbol{\omega}_{i}$ can be represented by a combination of performance measures of other states. This model with subsidy can be explained more intuitively through a two-arm system. In this system, the first arm is the original arm while the second arm (auxiliary arm) has only one state 0. In every slot we choose one of the two arms to activate. When the auxiliary arm is activated, we obtain a fixed reward $\lambda$. Our goal is to find a policy to maximize the long-term expected discounted reward by deciding which arm to choose in each slot. The objective function can thus be written as
$\max\limits_{\pi}\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}(\pi)+\lambda x_{0}^{1}(\pi),$
where $x_{0}^{1}(\pi)$ is the performance measure of state 0 being activated under policy $\pi$. Let $X$ be the set of all elements $(x_{0}^{1}(\pi),x_{\boldsymbol{\omega}}^{1}(\pi))_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}$ as $\pi$ traverses the admissible (feasible) policy set $\Pi$. We call $X$ the achievable region. Under this definition, the optimization objective function can be written as follows:
$\quad\max_{(x_{0}^{1},x_{\boldsymbol{\omega}}^{1})\in X}\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}+\lambda x_{0}^{1}.$ (OPT)
The core of solving this optimization problem is to mathematically characterize the achievable region $X$. The so-called conservation law refers to the fact that for any $(x_{0}^{1},x_{\boldsymbol{\omega}}^{1})\in X$, these components may satisfy certain equality or inequality constraints. For the above model, a trivial equality constraint is $x_{0}^{1}+\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega}}^{1}=\frac{1}{1-\beta}$. This is because the RMAB system has exactly one state in the active phase at each moment (time-conservation). In the following subsections, we will further explore the rich mathematical structure of performance measures for the two-arm system with countable states under the concept of achievable region.
2.5 Partial Conservation Law
In this section, we elaborate the relationship between performance measures for each state in the two-arm system. To measure the performance of the original arm under policy $\pi$, we define the following variables:
$T_{\boldsymbol{\omega}_{i}}^{\pi}$ $:=\mathbb{E}_{\pi}\left[\sum\limits_{t=0}^{+\infty}\beta^{t}\mathbbm{1}(a(t)=1)\bigg{|}\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}}),$ (22)
$R_{\boldsymbol{\omega}_{i}}^{\pi}$ $:=\mathbb{E}_{\pi}\left[\sum\limits_{t=0}^{+\infty}\beta^{t}R_{\boldsymbol{\omega}(t)}\mathbbm{1}(a(t)=1)\bigg{|}\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}}).$ (23)
Here $T_{\boldsymbol{\omega}_{i}}^{\pi}$ describes the expected total discounted time for the original arm to be activated in the two-arm system under policy $\pi$. We require that the policy $\pi$ only depends on the original arm. Similarly, $R_{\boldsymbol{\omega}_{i}}^{\Omega}$ represents the expected total discounted reward obtained from the original arm under policy $\pi$. Therefore, $T_{\boldsymbol{\omega}_{i}}^{\pi}$ and $R_{\boldsymbol{\omega}_{i}}^{\pi}$ also have the following expressions:
$T_{\boldsymbol{\omega}_{i}}^{\pi}$ $=\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega}_{0})}x_{\boldsymbol{\omega}}^{1}(\pi,\boldsymbol{\omega}_{i}),$ (24)
$R_{\boldsymbol{\omega}_{i}}^{\pi}$ $=\sum\limits_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega}_{0})}R_{\boldsymbol{\omega}}x_{\boldsymbol{\omega}}^{1}(\pi,\boldsymbol{\omega}_{i}).$ (25)
Clearly, the optimal policy partitions the belief state space $\Omega(\boldsymbol{\omega_{0}})$ of the original arm into two sets where the optimal actions should be active and passive respectively. For any $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, we call $\pi_{\Omega}$ an $\Omega$-priority policy if the original arm is activated when its current belief state falls into $\Omega$ and made passive when the state falls into $\Omega^{c}$. For $\Omega$-priority policy, we have
$T_{\boldsymbol{\omega}_{i}}^{\Omega}$ $=\mathbb{E}_{\pi_{\Omega}}\left[\sum\limits_{t=0}^{+\infty}\beta^{t}\mathbbm{1}(\boldsymbol{\omega}(t)\in\Omega)\bigg{|}\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}}),$ (26)
$R_{\boldsymbol{\omega}_{i}}^{\Omega}$ $=\mathbb{E}_{\pi_{\Omega}}\left[\sum\limits_{t=0}^{+\infty}\beta^{t}R_{\boldsymbol{\omega}(t)}\mathbbm{1}(\boldsymbol{\omega}(t)\in\Omega)\bigg{|}\boldsymbol{\omega}(0)=\boldsymbol{\omega}_{i}\right],\quad\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}}).$ (27)
From the above definition, we can easily give the following dynamic programming equations for these two variables:
$T_{\boldsymbol{\omega}_{i}}^{\Omega}=\begin{cases}1+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}T_{\boldsymbol{\omega}_{i}}^{\Omega},&\boldsymbol{\omega}_{i}\in\Omega\\ \beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}T_{\boldsymbol{\omega}_{i}}^{\Omega}&\boldsymbol{\omega}_{i}\notin\Omega\end{cases}$ (28)
and
$R_{\boldsymbol{\omega}_{i}}^{\Omega}=\begin{cases}R_{\boldsymbol{\omega}_{i}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}R_{\boldsymbol{\omega}_{i}}^{\Omega},&\boldsymbol{\omega}_{i}\in\Omega\\ \beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}R_{\boldsymbol{\omega}_{i}}^{\Omega}&\boldsymbol{\omega}_{i}\notin\Omega\end{cases}.$ (29)
If $\Omega(\boldsymbol{\omega_{0}})$ is a finite state space, we can directly solve for $T_{\boldsymbol{\omega_{i}}}$ and $R_{\boldsymbol{\omega_{i}}}$ from the two equation sets above. However, a direct solution is not available for countable state spaces. To further investigate the properties of these variables, we define $V_{\boldsymbol{\omega_{i}}}^{\pi}$ as the expected total discounted reward starting from state $\boldsymbol{\omega_{i}}$ under policy $\pi$. We have
$V_{\boldsymbol{\omega_{i}}}^{\pi}=R_{\boldsymbol{\omega_{i}}}^{\pi}+\lambda\left(\frac{1}{1-\beta}-T_{\boldsymbol{\omega_{i}}}^{\pi}\right).$ (30)
Let $V_{\boldsymbol{\omega_{i}},0}^{\pi}$ and $V_{\boldsymbol{\omega_{i}},1}^{\pi}$ be the expected total discounted rewards with the initial state $\boldsymbol{\omega_{i}}$ when taking the passive and active actions respectively, then
$V_{\boldsymbol{\omega_{i}},0}^{\pi}=\lambda+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}V_{\boldsymbol{\omega_{j}}}^{\pi},$ (31)
$V_{\boldsymbol{\omega_{i}},1}^{\pi}=R_{\boldsymbol{\omega_{i}}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}V_{\boldsymbol{\omega_{j}}}^{\pi}.$ (32)
If, under policy $\pi$, the expected total discounted reward obtained by active and passive actions are the same for the initial state $\boldsymbol{\omega_{i}}$, i.e., $V_{\boldsymbol{\omega_{i},0}}^{\pi}=V_{\boldsymbol{\omega_{i},1}}^{\pi}$, then we can solve for $\lambda$:
$\lambda=\frac{R_{\boldsymbol{\omega_{i}}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}(p_{ij}^{1}-p_{ij}^{0})R_{\boldsymbol{\omega_{j}}}^{\pi}}{1+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}(p_{ij}^{1}-p_{ij}^{0})T_{\boldsymbol{\omega_{j}}}^{\pi}}.$ (33)
Now define
$A_{\boldsymbol{\omega}_{i}}^{\Omega}$ $:=1+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}(p_{ij}^{1}-p_{ij}^{0})T_{\boldsymbol{\omega}_{j}}^{\Omega^{c}},$ (34)
$W_{\boldsymbol{\omega}_{i}}^{\Omega}$ $:=R_{\boldsymbol{\omega}_{i}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}(p_{ij}^{1}-p_{ij}^{0})R_{\boldsymbol{\omega}_{j}}^{\Omega^{c}}.$ (35)
For the auxiliary arm, we have $A_{0}^{\{0\}}=1$ and $W_{0}^{\{0\}}=\lambda$. Generally, we can extend these variables to the multi-arm case. Assume that there are two arms and their state spaces are $\Omega(\boldsymbol{\omega})$ and $\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ with $\Omega(\boldsymbol{\omega})\cap\Omega^{\prime}(\boldsymbol{\omega}^{\prime})=\emptyset$. Then for $\Omega\subset\Omega(\boldsymbol{\omega})$, $\Omega^{\prime}\subset\Omega^{\prime}(\boldsymbol{\omega}^{\prime})$ and $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega})$, we define $T_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=T_{\boldsymbol{\omega_{i}}}^{\Omega},R_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=R_{\boldsymbol{\omega_{i}}}^{\Omega},A_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=A_{\boldsymbol{\omega_{i}}}^{\Omega}$ and $W_{\boldsymbol{\omega_{i}}}^{\Omega\cup\Omega^{\prime}}:=W_{\boldsymbol{\omega_{i}}}^{\Omega}$. Previously, Niño-Mora provided the relationships among $T_{\boldsymbol{\omega_{i}}}^{\Omega}$, $R_{\boldsymbol{\omega_{i}}}^{\Omega}$, $A_{\boldsymbol{\omega_{i}}}^{\Omega}$, and $W_{\boldsymbol{\omega_{i}}}^{\Omega}$ for finite-state problems in (Niño-Mora 2001, 2002). Here, we prove that these conservation relations also hold for countable state spaces:
Theorem 2.4 (Decomposition Law)
Assume there exists a constant $C$, such that $|R_{\boldsymbol{\omega_{i}}}|\leq C$ holds for any $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ and $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, the following two identities hold:
$T_{\boldsymbol{\omega}_{j}}^{\pi}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega}_{j})$ $=T_{\boldsymbol{\omega}_{j}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega}_{j}),$ (36)
$R_{\boldsymbol{\omega}_{j}}^{\pi}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}W_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega}_{j})$ $=R_{\boldsymbol{\omega}_{j}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}W_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega}_{j}).$ (37)
Proof 2.5
Proof. Assume the initial belief state is $\boldsymbol{\omega_{k}}$. For convenience, abbreviate $x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{k}})$ and $x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{k}})$ as $x_{\boldsymbol{\omega_{i}}}^{1}$ and $x_{\boldsymbol{\omega_{i}}}^{0}$. In this problem, we also have the constraint equation:
$x_{\boldsymbol{\omega_{i}}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0}=e_{\boldsymbol{\omega_{i}}}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}+\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0},\quad\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}}).$
Since $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{1}+x_{\boldsymbol{\omega_{i}}}^{0})=\frac{1}{1-\beta}$, and for any $\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})$, $x_{\boldsymbol{\omega_{i}}}^{1},x_{\boldsymbol{\omega_{i}}}^{0}\geq 0$, thus by the Cauchy criterion for series convergence, $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}$ and $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{0}$ are both absolutely convergent. Due to the countability of $\Omega(\boldsymbol{\omega_{0}})$, summing the above constraint equations gives:
$\frac{1}{1-\beta}=\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0}.$ (38)
Therefore, $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}$ and $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0}$ are both convergent and their summation orders are exchangeable. By constraint equations and convergence of series, we have
$\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{1}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1})T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{0}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0})T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}=T_{\boldsymbol{\omega_{k}}}^{\Omega^{c}}.$ (39)
Since $T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}$ has the upper bound $\frac{1}{1-\beta}$, the series in the above equation converge. In this equation,
$\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{1}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1})T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{1}x_{\boldsymbol{\omega_{j}}}^{1}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}x_{\boldsymbol{\omega_{i}}}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}(T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}})$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}})+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}x_{\boldsymbol{\omega_{i}}}^{1}(T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{1}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}})$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(1-A_{\boldsymbol{\omega_{i}}}^{\Omega})+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}x_{\boldsymbol{\omega_{i}}}^{1}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}-\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}x_{\boldsymbol{\omega_{i}}}^{1}A_{\boldsymbol{\omega_{i}}}^{\Omega}.$
Here, the third equality exchanges the order of summation and the sixth equality uses the equations of $A_{\boldsymbol{\omega_{i}}}^{\Omega}$ and $W_{\boldsymbol{\omega_{i}}}^{\Omega}$. Similarly,
$\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}(x_{\boldsymbol{\omega_{i}}}^{0}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ji}^{0}x_{\boldsymbol{\omega_{j}}}^{0})T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}})+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}x_{\boldsymbol{\omega_{i}}}^{0}(T_{\boldsymbol{\omega_{i}}}^{\Omega^{c}}-\beta\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}p_{ij}^{0}T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}})$
$=$ $\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}x_{\boldsymbol{\omega_{i}}}^{0}A_{\boldsymbol{\omega_{i}}}^{\Omega}.$
Note that
$\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{k}})=T_{\boldsymbol{\omega_{k}}}^{\pi}.$ (40)
We obtain
$T_{\boldsymbol{\omega_{k}}}^{\pi}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{k}})=T_{\boldsymbol{\omega_{k}}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{k}}).$ (41)
Similarly we have
$R_{\boldsymbol{\omega_{k}}}^{\pi}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}W_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{k}})=R_{\boldsymbol{\omega_{k}}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}W_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{k}}).$ (42)
[…]
We can also perform certain deformations on two equations. Note that $x_{0}^{1}(\pi,\boldsymbol{\omega_{j}})=\frac{1}{1-\beta}-T_{\boldsymbol{\omega_{j}}}^{\pi}$, equation (36) can be written as
$x_{0}^{1}(\pi,\boldsymbol{\omega_{j}})+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{j}})=\frac{1}{1-\beta}-T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{j}}).$ (43)
Similarly, we can convert (37) to
$\lambda x_{0}^{1}(\pi,\boldsymbol{\omega_{j}})+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}W_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{j}})=V_{\boldsymbol{\omega_{j}}}^{\pi}-R_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega^{c}}W_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{0}(\pi,\boldsymbol{\omega_{j}}).$ (44)
If for any $\boldsymbol{\omega_{i}}\in\Omega^{c}$ we have $A_{\boldsymbol{\omega_{i}}}^{\Omega}\geq 0$, then
$x_{0}^{1}(\pi,\boldsymbol{\omega_{j}})+\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega}A_{\boldsymbol{\omega_{i}}}^{\Omega}x_{\boldsymbol{\omega_{i}}}^{1}(\pi,\boldsymbol{\omega_{j}})\geq\frac{1}{1-\beta}-T_{\boldsymbol{\omega_{j}}}^{\Omega^{c}}.$
The equality holds when policy $\pi$ gives higher priority to state 0 (the second arm) than states in $\Omega^{c}$. However, not all $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$ have such property, so we need to consider a family of sets $\tilde{\Omega}\subset 2^{\Omega(\boldsymbol{\omega_{0}})}$, which satisfies $\Omega(\boldsymbol{\omega_{0}})\in\tilde{\Omega}$. For any $\Omega\in\tilde{\Omega}$, define $b(\Omega):=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\},b(\{0\}\cup\Omega)=\inf\limits_{\pi}\{x_{0}^{1}(\pi)+\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$. Here we give the following definition of partial conservation law generalized from Niño-Mora (2001):
Definition 2.6 (Partial Conservation Law (PCL))
The problem is said to satisfy the partial conservation law with respect to $\tilde{\Omega}$ if, for any $\Omega\subset\tilde{\Omega}$ and policy $\pi$, the following hold:
$x_{0}^{1}(\pi)+\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\geq b(\{0\}\cup\Omega),$
$\sum\limits_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\geq b(\Omega),$
$x_{0}^{1}(\pi)+\sum\limits_{\boldsymbol{\omega}\in\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)=b(\{0\}\cup\Omega(\boldsymbol{\omega}_{0})).$
The first inequality holds with equality if policy $\pi$ gives priority to 0 over $\Omega$, and the second inequality holds with equality if policy $\pi$ gives priority to $\Omega^{c}$ over 0.
Conservation law limits the range of feasible region composed of performance measures of various states through a series of constraints. However, we are not yet ready to judge whether the objective function achieves the optimal value in this constrained region. In order to further characterize the indices of priority policies, we need to restrict the aforementioned $\tilde{\Omega}$. We first give the following definition on a chain of subsets of $\tilde{\Omega}$:
Definition 2.7 (Full subset chain of $\tilde{\Omega}$)
We call collection family $C=\{\Omega_{\boldsymbol{\omega_{i}}}^{C}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\subset\tilde{\Omega}$ is a full subset chain of $\tilde{\Omega}$, if it satisfies the following conditions:
(i) $\forall\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}}),\boldsymbol{\omega_{i}}\in\Omega_{\boldsymbol{\omega_{i}}}^{C}$;
(ii) $\forall\boldsymbol{\omega_{i}},\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}}),\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$, there must be $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$ or $\Omega_{\boldsymbol{\omega_{j}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}^{C}$, and if $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, then $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}^{C}$.
Furthermore, the entire full subset chain of $\tilde{\Omega}$ is denoted as $\mathcal{C}(\tilde{\Omega})$.
Now assume $\tilde{\Omega}$ satisfies $\mathcal{C}(\tilde{\Omega})\neq\emptyset$. For $C\in\mathcal{C}(\tilde{\Omega})$, we can use the equation (33) to formally construct variable $\lambda_{\boldsymbol{\omega_{i}}}^{C}=\frac{W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{C}}}{A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{C}}}$. With the above concept, we provide the definition of PCL-indexability for countable state problems as follows:
Definition 2.8 (PCL-indexability)
We call the problem is PCL-indexable w.r.t. $\tilde{\Omega}$ if
(i) The problem satisfies PCL w.r.t. $\tilde{\Omega}$;
(ii) $\exists!C\in\mathcal{C}(\tilde{\Omega})$, s.t.$\forall\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$ satisfying $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, there is $\lambda_{\boldsymbol{\omega_{i}}}^{C}<\lambda_{\boldsymbol{\omega_{j}}}^{C}$.
Condition (i) in Definition 2.8 characterizes part of the boundary of the achievable region through the constraints between performance measures for an infinite dimensional linear programming problem, and condition (ii) means that the objective function can take its maximum value at a unique vertex in the achievable region while the policy corresponding to this vertex is a state priority policy obtained by sorting according to the size of the index $\lambda_{\boldsymbol{\omega}}$. Next, we show that if the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then we can construct a unique full subset chain of $\tilde{\Omega}_{0}:=\tilde{\Omega}\cup\{\{0\}\cup\Omega,\Omega\in\tilde{\Omega}\}$. Assume $\exists!C\in\mathcal{C}(\tilde{\Omega})$, s.t.$\forall\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$ with $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, we have $\lambda_{\boldsymbol{\omega_{i}}}^{C}<\lambda_{\boldsymbol{\omega_{j}}}^{C}$. We can construct the full subset chain $C^{\prime}$ of $\tilde{\Omega}_{0}$ as follows:
$\Omega_{\boldsymbol{\omega}}^{C^{\prime}}=\begin{cases}\Omega_{\boldsymbol{\omega}}^{C}\cup\{0\},&\lambda_{\boldsymbol{\omega}}^{C}>\lambda\\ \Omega_{\boldsymbol{\omega}}^{C},&\lambda_{\boldsymbol{\omega}}^{C}\leq\lambda\end{cases}$
and
$\Omega_{0}^{C^{\prime}}=\{0\}\cup\bigcup\limits_{\boldsymbol{\omega}:\lambda_{\boldsymbol{\omega}}\leq\lambda}\Omega_{\boldsymbol{\omega}}^{C}.$
Without loss of generality, here we can assume $\bigcup\limits_{\boldsymbol{\omega}:\lambda_{\boldsymbol{\omega}}\leq\lambda}\Omega_{\boldsymbol{\omega}}^{C}\in\tilde{\Omega}$, otherwise we just need to add it to $\tilde{\Omega}$. Under such a construction, it is easy to verify that $C^{\prime}$ is a full subset chain of $\tilde{\Omega}_{0}$, and it satisfies the second condition of Definition 2.8. Furthermore, we have the following characterization of the unique full subset chain of $\tilde{\Omega}_{0}$:
Lemma 2.9
Assume the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then for the unique full subset chain $C^{\prime}$ constructed above, it has the following index characterization:
$\Omega_{\boldsymbol{\omega_{i}}}=\{\boldsymbol{\omega_{k}}\in\Omega(\boldsymbol{\omega_{0}})\cup\{0\}:\lambda_{\boldsymbol{\omega_{k}}}\leq\lambda_{\boldsymbol{\omega_{i}}}\}.$ (45)
Proof 2.10
[…]
By combining Theorem 2.4 and Lemma 2.9, the following relationship can be established:
Lemma 2.11
$(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}})\frac{W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}}{A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}}$.
Proof 2.12
Proof. First we consider the case that $\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$. Let $\Omega=\Omega_{\boldsymbol{\omega_{i}}},\pi=\pi_{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}$ in Theorem 2.4, then
$T_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}$ $=T_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}}+A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}x_{\boldsymbol{\omega_{i}}}^{1}(\pi_{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}},\boldsymbol{\omega_{j}}),$ (46)
$R_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}$ $=R_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}}+W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}x_{\boldsymbol{\omega_{i}}}^{1}(\pi_{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}},\boldsymbol{\omega_{j}}).$ (47)
Thus we have
$W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}(T_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}-T_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}})=A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}(R_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}-R_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}}).$ (48)
Since the above equation holds for any $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, we have
$W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}(p_{ij}^{1}-p_{ij}^{0})(T_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}-T_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}})=A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}\sum\limits_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}(p_{ij}^{1}-p_{ij}^{0})(R_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}\cup\{\boldsymbol{\omega_{i}}\}}-R_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{c}}).$
From the dynamic programming equations we obtain
$W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}})=A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}(W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}}).$
[…]
2.6 Optimality of Index Policy
Now we address the main issue whether the sequence of $\{\lambda_{\boldsymbol{\omega}}\}$-priorities when the problem satisfies PCL-indexability achieves optimality. Since the problem satisfies PCL, the partial conservation law inequality describes a region for an infinite-dimensional linear programming problem. First, we give a brief introduction to the basic theory of infinite-dimensional linear programming. Let $X$ and $Y$ be real linear vector spaces and the corresponding positive cones are $P_{X}$ and $P_{Y}$. Let $X$ be partially ordered by relation $\geq$, defined by $x\geq y$ if $x-y\in P_{X}$. Let $X^{*}$ be the dual space of $X$. For any $x^{*}\in X^{*}$, $x^{*}$ is a linear functional on $X$ and the image of $x\in X$ is denoted by $\langle x,x^{*}\rangle$. Similarly, space $Y$ for the dual problem also derives $Y^{*}$ and $P_{Y^{*}}$. Let $A$ be a linear map from $X$ to $Y$, $y\in Y$ and $x^{*}\in X^{*}$. The primal linear problem is
$\max_{x}\langle x,x^{*}\rangle$ (LP)
$s.t.$ $Ax-y\in P_{Y}$
$x\in P_{X}$
Let $A^{*}$ be the dual linear map from $Y^{*}$ to $X^{*}$, then the dual linear program can be written as:
$\min_{y^{*}}\langle y,y^{*}\rangle$ (LD)
$s.t.$ $A^{*}y^{*}-x^{*}\in P_{X^{*}}$
$-y^{*}\in P_{Y^{*}}$
Similar to the case of finite dimensions, weak duality relation can be established between (LP) and (LD).
Theorem 2.13
(Anderson and Nash 1987) Assume $x$ is feasible for (LP) and $y^{*}$ is feasible for (LD), then
$\langle y,y^{*}\rangle\geq\langle x,x^{*}\rangle.$ (49)
Equality holds if and only if
$\langle y-Ax,y^{*}\rangle=0,\quad\langle x,A^{*}y^{*}-x^{*}\rangle=0$ (50)
in which case we say that $x,y^{*}$ are complementary slack. If $x,y^{*}$ are feasible and complementary slack then $x$ and $y^{*}$ are optimal solutions to (LP) and (LD) respectively.
Next, we aim to prove that if the problem satisfies PCL-indexability w.r.t. $\tilde{\Omega}$, then the problem is also Whittle indexable, i.e., PCL-indexability is a sufficient condition for Whittle indexability. First, we obtain the optimality of the $\{\lambda_{\boldsymbol{\omega}}\}$-priority policy in the two-arm system. Assume the problem is PCL-indexable w.r.t. $\tilde{\Omega}$. For simplicity, we now denote the joint state space $\{0\}\cup\Omega(\boldsymbol{\omega_{0}})$ by $\Omega(\boldsymbol{\omega_{0}})$, i.e. we view state 0 as a part of the whole two-arm system’s state space and denote $\tilde{\Omega}\cup\{\{0\}\cup\Omega:\Omega\in\tilde{\Omega}\}$ by $\tilde{\Omega}$. Consider the space $X=\{\{x_{\boldsymbol{\omega_{i}}}^{1}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}|\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}|x_{\boldsymbol{\omega_{i}}}^{1}|<+\infty\}$. For any reward vector $\boldsymbol{R}=(R_{\boldsymbol{\omega_{i}}})_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$, define the inner product $\langle\boldsymbol{R},x^{1}\rangle:=\sum\limits_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{i}}}^{1}R_{\boldsymbol{\omega_{i}}}$. Thus $\boldsymbol{R}\in X^{*}$. Furthermore, consider the variable space $Y$ of the dual problem, where $Y$ is a space of set functions on $\tilde{\Omega}$. Let $b\in Y$, for $\Omega\in\tilde{\Omega}$, define $b(\Omega)$ as
$b(\Omega):=\inf\limits_{\pi}\{\sum_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi):\pi\in\Pi\}.$
For $A:X\rightarrow Y,~{}x^{1}\mapsto Ax^{1}$, define $Ax^{1}(\Omega)=\sum_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1},\Omega\in\tilde{\Omega}$. Let $Y$ be the linear vector space spanned by $b$ and $Ax^{1},~{}x^{1}\in X$. The positive cone is $P_{Y}=\{y\in Y:y(\Omega)\geq 0,\Omega\subset\Omega(\boldsymbol{\omega_{0}}),y(\Omega(\boldsymbol{\omega_{0}}))=0\}$. By Lemma 2.9, $\Omega_{\boldsymbol{\omega_{i}}}=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}\leq\lambda_{\boldsymbol{\omega_{i}}}\}$. Let $\Omega_{\boldsymbol{\omega_{i}}}^{-}:=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}\}$. In infinite dimensional problems, for any $y\in Y$, define $\lambda^{*}\in Y^{*}$ as follows:
$\langle y,\lambda^{*}\rangle:=\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}\left[y(\Omega_{\boldsymbol{\omega_{i}}})-y(\Omega_{\boldsymbol{\omega_{i}}}^{-})\right].$
Let $x^{W}\in X$ be the performance measure given by the $\{\lambda_{\boldsymbol{\omega_{i}}}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$-priority policy. Our goal is to prove that $x^{W}$ and $\lambda^{*}$ are optimal solutions to (LP) and (LD), respectively. Due to the assumption that the problem satisfies PCL-indexability w.r.t. $\tilde{\Omega}$, it satisfies the PCL condition and $x^{W}$ is a feasible solution of (LP). By the weak duality principle, we need to prove the following two statements:
(i) $\lambda^{*}$ is a feasible solution to (LD);
(ii) $\langle Ax^{W}-b,\lambda^{*}\rangle=0$ and $\langle x^{W},A^{*}\lambda^{*}-\boldsymbol{R}\rangle=0$.
For the first complementary slackness condition in (ii), by the definition of $x^{W}$ and $\lambda^{*}$, we have
$\langle Ax^{W},\lambda^{*}\rangle$ $=\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}(Ax^{W}(\Omega_{\boldsymbol{\omega_{i}}})-Ax^{W}(\Omega_{\boldsymbol{\omega_{i}}}^{-}))$
$=\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}\left[b(\Omega_{\boldsymbol{\omega_{i}}})-b(\Omega_{\boldsymbol{\omega_{i}}}^{-})\right]$
$=\langle b,\lambda^{*}\rangle.$
Thus $\langle Ax^{W}-b,\lambda^{*}\rangle=0$ holds. The establishment of the second complementary slackness condition requires an important equation that is trivially satisfied in finite-state problems. However, for the case of countable states considered here, the situation is much more complex and we prove it in the theorem below based on the following assumption. {assumption} Sequence $\{\lambda_{\boldsymbol{\omega}}\}_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})\cup\{0\}}$ has finitely many limit points. In fact, since state update rules are themselves finite and the consideration of stationary policies is often sufficient for long-term optimization problems, the above assumption is naturally satisfied in practice. Based on this assumption, we give the following important equation.
Theorem 2.14
$R_{\boldsymbol{\omega_{j}}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{j}}}}+\sum\limits_{\boldsymbol{\omega_{i}}:\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}}(W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}})$.
In order to prove this theorem, we first establish an easy lemma. For simplicity, we denote the state space $\Omega_{\boldsymbol{\omega_{0}}}\cup\{0\}$ by $\mathcal{M}$ and state $\boldsymbol{\omega_{i}}$ by $i$.
Lemma 2.15
For $S\subset\mathcal{M}$, let $R_{T,i}^{S}$ denote the expected discounted reward obtained from initial belief state $i$ under the $S$-priority policy over the first $T$ time steps. Then we have $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{S}=R_{i}^{S}$. In fact, this limit process is independent of $j$, i.e. $R_{T,i}^{S}\rightrightarrows R_{i}^{S}(T\rightarrow+\infty)$.
Proof 2.16
Proof. Let $p_{T,ij}^{S}$ be the probability of being in state $j$ after $T$ time steps starting from state $i$ under the $S$-priority policy. Then we have
$R_{i}^{S}=R_{T,i}^{S}+\beta^{T}\sum\limits_{j\in\mathcal{M}}p_{T,ij}^{S}R_{j}^{S}.$
[…]
To prove Theorem 2.14, we only need to consider the case where the sequence $\{\lambda_{i}\}_{i\in\mathcal{M}}$ has one limit point. For notation convenience, we renumber the states in $\mathcal{M}$ as $\{*,1,2,\cdots\}$. The state indices in the sequence have the following relationship:
[…]
In this figure, $\lambda_{*}$ is the unique limit point of $\{\lambda_{i}\}_{i\in\mathcal{M}}$ and $\{\lambda_{2n-1}\}\searrow\lambda_{*}$ while $\{\lambda_{2n}\}\nearrow\lambda_{*}$. Recall that $\Omega_{i}=\{j\in\mathcal{M}:\lambda_{j}\leq\lambda_{i}\}$, we have the following lemma:
Lemma 2.17
$\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}},~{}\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n}^{c}}=R_{i}^{\Omega_{*}^{-c}},~{}\forall i\in\mathcal{M}$.
Proof 2.18
Proof. We first prove that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}}$. We will use induction to show that for any $T$, $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$. When $T=1$, if $i\notin\Omega_{*}^{c}$, then for any $n$, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=0$; If $i\in\Omega_{*}^{c}$, since $\lambda_{*}$ is the limit point, there $\exists N$ s.t. $i\in\Omega_{2n-1}^{c}$ when $n>N$. In this case, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=R_{i}$. Thus $\lim\limits_{n\rightarrow+\infty}R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}$ holds for every $i\in\mathcal{M}$.
When $T\geq 1$, we assume $\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$ holds for every $i\in\mathcal{M}$, then we need to consider the following two situations:
(i) If $i\notin\Omega_{*}^{c}$, we have
$R_{T+1,i}^{\Omega_{2n-1}^{c}}=\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{0}R_{T,j}^{\Omega_{2n-1}^{c}},R_{T+1,i}^{\Omega_{*}^{c}}=\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{0}R_{T,j}^{\Omega_{*}^{c}}.$
Thus
$|R_{T+1,i}^{\Omega_{2n-1}^{c}}-R_{T+1,i}^{\Omega_{*}^{c}}|\leq\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{0}|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|.$
For any $\varepsilon>0$, since $p_{i*}^{0}+\sum\limits_{j=1}^{+\infty}p_{ij}^{0}=1$, there $\exists J$ s.t. $\sum\limits_{j=J+1}^{+\infty}p_{ij}^{0}\leq\frac{\varepsilon(1-\beta)}{4\beta C}$. For $j=*,1,\cdots,J$, by the induction hypothesis, $\exists N(j)$ s.t. when $n>N(j)$, $|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|\leq\frac{\varepsilon}{2\beta}$. Let $N=\max\limits_{j\in\{*,1,\cdots,J\}}N(j)$, then when $n>N$, we have
$|R_{T+1,i}^{\Omega_{2n-1}^{c}}-R_{T+1,i}^{\Omega_{*}^{c}}|$ $\leq\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{0}|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|$
$=\beta\sum\limits_{j\in\{*,1,\cdots,J\}}p_{ij}^{0}|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|+\beta\sum\limits_{j=J+1}^{+\infty}p_{ij}^{0}|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|$
$\leq\beta\cdot\left(\sum\limits_{j\in\{*,1,\cdots,J\}}p_{ij}^{0}\right)\cdot\frac{\varepsilon}{2\beta}+\beta\cdot\left(\sum\limits_{j=J+1}^{+\infty}p_{ij}^{0}\right)\cdot\frac{2C}{1-\beta}$
$\leq\beta\cdot\frac{\varepsilon}{2\beta}+\beta\cdot 2C\cdot\frac{\varepsilon}{4\beta C}=\varepsilon.$
Therefore, $\lim\limits_{n\rightarrow+\infty}R_{T+1,i}^{\Omega_{2n-1}^{c}}=R_{T+1,i}^{\Omega_{*}^{c}}$.
(ii) If $i\in\Omega_{*}^{c}$, then there $\exists N$ s.t. $i\in\Omega_{n}^{c}$ when $n>N$. In this case, we have
$R_{T+1,i}^{\Omega_{2n-1}^{c}}=R_{i}+\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{1}R_{T,j}^{\Omega_{2n-1}^{c}},R_{T+1,i}^{\Omega_{*}^{c}}=R_{i}+\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{1}R_{T,j}^{\Omega_{*}^{c}}.$
Thus
$|R_{T+1,i}^{\Omega_{2n-1}^{c}}-R_{T+1,i}^{\Omega_{*}^{c}}|\leq\beta\sum\limits_{j\in\mathcal{M}}p_{ij}^{1}|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|.$
Using the discussion similar to (i) one can derive that $\lim\limits_{n\rightarrow+\infty}R_{T+1,i}^{\Omega_{2n-1}^{c}}=R_{T+1,i}^{\Omega_{*}^{c}}$. By induction, we know that for any $T$, $\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$.
Next we will prove that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}}$. $\forall\varepsilon>0$, let $T=\lfloor\frac{\log\varepsilon+\log(1-\beta)-\log C}{\log\beta}\rfloor$, then when $T_{2}>T_{1}>T$, for every $n$, we have
$R_{T_{2},i}^{\Omega_{2n-1}^{c}}=R_{T_{1},i}^{\Omega_{2n-1}^{c}}+\beta^{T_{1}}\sum\limits_{j\in\mathcal{M}}p_{T_{1},ij}^{\Omega_{2n-1}^{c}}R_{T_{2}-T_{1},j}^{\Omega_{2n-1}^{c}}.$
Thus
$|R_{T_{2},i}^{\Omega_{2n-1}^{c}}-R_{T_{1},i}^{\Omega_{2n-1}^{c}}|$ $=\beta^{T_{1}}|\sum\limits_{j\in\mathcal{M}}p_{T_{1},ij}^{\Omega_{2n-1}^{c}}R_{T_{2}-T_{1},j}^{\Omega_{2n-1}^{c}}|$
$\leq\frac{\beta^{T}C}{1-\beta}<\varepsilon.$
By Theorem 7.11 in Rudin et al. (1976), $\lim\limits_{n\rightarrow+\infty}\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=\lim\limits_{T\rightarrow+\infty}\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$. By Lemma 2.15,
$\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}$ $=\lim\limits_{n\rightarrow+\infty}\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$
$=\lim\limits_{T\rightarrow+\infty}\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$
$=\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{*}^{c}}$
$=R_{i}^{\Omega_{*}^{c}},\quad\forall i\in\mathcal{M}.$
[…]
Lemma 2.19
$\forall i\in\mathcal{M},~{}\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}},~{}\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n}}=W_{i}^{\Omega_{*}^{-}},~{}\lim\limits_{n\rightarrow+\infty}A_{i}^{\Omega_{2n+1}}=A_{i}^{\Omega_{*}}$, and $\lim\limits_{n\rightarrow+\infty}A_{i}^{\Omega_{2n}}=A_{i}^{\Omega_{*}^{-}}$.
Proof 2.20
Proof. Here we only prove that $\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}}$, other identities can be proved similarly. By definition,
$W_{i}^{\Omega_{2n+1}}=R_{i}+\beta\sum\limits_{j\in\mathcal{M}}(p_{ij}^{1}-p_{ij}^{0})R_{j}^{\Omega_{2n+1}^{c}},W_{i}^{\Omega_{*}}=R_{*}+\beta\sum\limits_{j\in\mathcal{M}}(p_{ij}^{1}-p_{ij}^{0})R_{j}^{\Omega_{*}^{c}}.$
$\forall\varepsilon>0$, there $\exists J$, s.t. $\sum\limits_{j=J+1}^{+\infty}(p_{ij}^{1}+p_{ij}^{0})\leq\frac{\varepsilon(1-\beta)}{4\beta C}$. For $j=*,1,\cdots,J$, since $\lim\limits_{n\rightarrow+\infty}R_{j}^{\Omega_{2n+1}^{c}}=R_{j}^{\Omega_{*}^{c}}$, there $\exists N(j)$ s.t. when $n>N(j)$, $|R_{j}^{\Omega_{2n+1}^{c}}-R_{j}^{\Omega_{*}^{c}}|<\frac{\varepsilon}{4\beta}$. Let $N=\max\limits_{j\in\{*,1,\cdots,J\}}N(j)$, then when $n>N$, we have
$|W_{i}^{\Omega_{2n+1}}-W_{i}^{\Omega_{*}}|$ $=|\beta\sum\limits_{j\in\mathcal{M}}(p_{ij}^{1}-p_{ij}^{0})(R_{j}^{\Omega_{2n+1}^{c}}-R_{j}^{\Omega_{*}^{c}})|$
$\leq\beta\sum\limits_{j\in\mathcal{M}}|p_{ij}^{1}-p_{ij}^{0}||R_{j}^{\Omega_{2n+1}^{c}}-R_{j}^{\Omega_{*}^{c}}|$
$\leq\beta\sum\limits_{j\in\{*,1,\cdots,J\}}(p_{ij}^{1}+p_{ij}^{0})|R_{j}^{\Omega_{2n+1}^{c}}-R_{j}^{\Omega_{*}^{c}}|+\beta\sum\limits_{j=J+1}^{+\infty}(p_{ij}^{1}+p_{ij}^{0})|R_{j}^{\Omega_{2n+1}^{c}}-R_{j}^{\Omega_{*}^{c}}|$
$\leq\beta\cdot\frac{\varepsilon}{4\beta}\cdot\sum\limits_{j\in\{*,1,\cdots,J\}}(p_{ij}^{1}+p_{ij}^{0})+\beta\cdot\frac{2C}{1-\beta}\cdot\sum\limits_{j=J+1}^{+\infty}(p_{ij}^{1}+p_{ij}^{0})$
$\leq\frac{\varepsilon}{4}\cdot 2+\frac{2\beta C}{1\beta}\cdot\frac{\varepsilon(1-\beta)}{4\beta C}$
$=\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.$
[…]
Proof 2.21
Proof of Theorem 2.14. We will separately consider three cases where $j$ is odd, $j=*$ and $j$ is even.
(i) If $j$ is odd, we can assume $j=2N-1$, then
$W_{j}^{\Omega_{j}}+\sum\limits_{i:\lambda_{i}>\lambda_{j}}(W_{j}^{\Omega_{i}}-W_{j}^{\Omega_{i}^{-}})$ $=W_{2N-1}^{\Omega_{2N-1}}+\sum\limits_{n=1}^{N}(W_{2N-1}^{\Omega_{2n-1}}-W_{2N-1}^{\Omega_{2n+1}})$
$=W_{2N-1}^{\Omega_{1}}=R_{2N-1}.$
(ii) If $j=*$, then by Lemma 2.19,
$W_{*}^{\Omega_{*}}+\sum\limits_{i:\lambda_{i}>\lambda_{*}}(W_{*}^{\Omega_{i}}-W_{*}^{\Omega_{i}^{-}})$ $=W_{*}^{\Omega_{*}}+\sum\limits_{n=1}^{+\infty}(W_{*}^{\Omega_{2n-1}}-W_{*}^{\Omega_{2n+1}})$
$=W_{*}^{\Omega_{*}}+\lim\limits_{N\rightarrow+\infty}\sum\limits_{n=1}^{N}(W_{*}^{\Omega_{2n-1}}-W_{*}^{\Omega_{2n+1}})$
$=W_{*}^{\Omega_{*}}+W_{*}^{\Omega_{1}}-\lim\limits_{N\rightarrow+\infty}W_{*}^{\Omega_{2N+1}}$
$=R_{*}.$
(iii) If $j$ is even, we can assume $j=2N$, then by Lemma 2.19, we have:
[…]
$\sum\limits_{i:\lambda_{i}>\lambda_{j}}(W_{j}^{\Omega_{i}}-W_{j}^{\Omega_{i}^{-}})$ $=\sum\limits_{n=1}^{+\infty}(W_{j}^{\Omega_{2n-1}}-W_{j}^{\Omega_{2n-1}^{-}})+\sum\limits_{n=N+1}^{+\infty}(W_{j}^{\Omega_{2n}}-W_{j}^{\Omega_{2n}^{-}})+W_{j}^{\Omega_{*}}-W_{j}^{\Omega_{*}^{-}}$
$=\sum\limits_{n=1}^{+\infty}(W_{j}^{\Omega_{2n-1}}-W_{j}^{\Omega_{2n+1}})+\sum\limits_{n=N+1}^{+\infty}(W_{j}^{\Omega_{2n}}-W_{j}^{\Omega_{2n-2}})+W_{j}^{\Omega_{*}}-W_{j}^{\Omega_{*}^{-}}$
$=W_{j}^{\Omega_{1}}-\lim\limits_{n\rightarrow+\infty}W_{j}^{\Omega_{2n+1}}-W_{j}^{\Omega_{j}}+\lim\limits_{n\rightarrow+\infty}W_{j}^{\Omega_{2n}}+W_{j}^{\Omega_{*}}-W_{j}^{\Omega_{*}^{-}}$
$=R_{j}-W_{j}^{\Omega_{j}}.$
[…]
Theorem 2.14 is a natural generalization of the key equation in finite-state problems to countable-state problems. Based on Theorem 2.14, we can establish the second relaxed complementary condition:
Lemma 2.22
$\langle x^{W},A^{*}\lambda^{*}-R\rangle=0$.
Proof 2.23
Proof. By Lemma 2.11 and Theorem 2.14, we have
$A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{j}}}}\lambda_{\boldsymbol{\omega_{j}}}$ $=W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{j}}}}$
$=R_{\boldsymbol{\omega_{j}}}-\sum\limits_{\boldsymbol{\omega_{i}}:\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}}(W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}})$
$=R_{\boldsymbol{\omega_{j}}}-\sum\limits_{\boldsymbol{\omega_{i}}:\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}}(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}})\lambda_{\boldsymbol{\omega_{i}}}.$
$\forall x^{1}\in X$, following from Proposition 4.3 in Frostig and Weiss (1999), we have
$\langle x^{1},A^{*}\lambda^{*}\rangle=\langle Ax^{1},\lambda^{*}\rangle$
$=$ $\sum_{\boldsymbol{\omega_{i}}\in\Omega(\omega_{0})}\lambda_{\boldsymbol{\omega_{i}}}\left[Ax^{1}(\Omega_{\boldsymbol{\omega_{i}}})-Ax^{1}(\Omega_{\boldsymbol{\omega_{i}}}^{-})\right]$
$=$ $\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}\left[\sum_{\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{i}}}}A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}x^{1}_{\boldsymbol{\omega_{j}}}-\sum_{\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{i}}}^{-}}A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}}x_{\boldsymbol{\omega_{j}}}^{1}\right]$
$=$ $\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}x_{\boldsymbol{\omega_{i}}}^{1}-\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}\sum_{\boldsymbol{\omega_{j}}:\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}}(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}})x_{\boldsymbol{\omega_{j}}}^{1}$
$=$ $\sum_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}\lambda_{\boldsymbol{\omega_{i}}}A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}x_{\boldsymbol{\omega_{i}}}^{1}-\sum_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{j}}}^{1}\sum_{\boldsymbol{\omega_{i}}:\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}}\lambda_{\boldsymbol{\omega_{i}}}(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}})$
$=$ $\sum_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{j}}}^{1}\left[\lambda_{\boldsymbol{\omega_{j}}}A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{j}}}}-\sum_{\boldsymbol{\omega_{i}}:\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}}\lambda_{\boldsymbol{\omega_{i}}}(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{-}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}})\right]$
$=$ $\sum_{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})}x_{\boldsymbol{\omega_{j}}}^{1}R_{\boldsymbol{\omega_{j}}}$
$=$ $\langle x^{1},R\rangle.$
Thus $\langle x^{W},A^{*}\lambda^{*}-R\rangle=0$Halmos
From the proof of Lemma 2.22, we know that $\lambda^{*}$ satisfies $A^{*}\lambda^{*}-\boldsymbol{R}\in P_{X^{*}}$. Therefore, the remaining condition in the weak duality theorem is $-\lambda^{*}\in P_{Y^{*}}$. To prove this statement, we need to show that, $\forall y\in P_{Y},\langle y,~{}\lambda^{*}\rangle\leq 0$. In fact,
$\langle y,\lambda^{*}\rangle$
$=$ $\sum\limits_{i\in\mathcal{M}}\lambda_{i}\left[y(\Omega_{i})-y(\Omega_{i}^{-})\right]$
$=$ $\sum\limits_{n=1}^{+\infty}\lambda_{2n-1}\left[y(\Omega_{2n-1})-y(\Omega_{2n-1}^{-})\right]+\lambda_{*}\left[y(\Omega_{*})-y(\Omega_{*}^{-})\right]+\sum\limits_{n=1}^{+\infty}\lambda_{2n}\left[y(\Omega_{2n})-y(\Omega_{2n}^{-})\right]$
$=$ $\sum\limits_{n=1}^{+\infty}\lambda_{2n-1}\left[y(\Omega_{2n-1})-y(\Omega_{2n+1})\right]+\lambda_{*}\left[y(\Omega_{*})-y(\Omega_{*}^{-})\right]+\sum\limits_{n=2}^{+\infty}\lambda_{2n}\left[y(\Omega_{2n})-y(\Omega_{2n-2})\right]$
$+\lambda_{2}y(\Omega_{2})$
$=$ $\lambda_{1}y(\Omega_{1})+\sum\limits_{n=2}^{+\infty}(\lambda_{2n-1}-\lambda_{2n-3})y(\lambda_{2n-1})-\lim\limits_{n\rightarrow+\infty}\lambda_{2n-1}y(\Omega_{2n+1})+\lambda_{*}y(\Omega_{*})-\lambda_{*}y(\Omega_{*}^{-})$
$+\sum\limits_{n=1}^{+\infty}(\lambda_{2n}-\lambda_{2n+2})y(\Omega_{2n})+\lim\limits_{n\rightarrow+\infty}\lambda_{2n+2}y(\Omega_{2n+2}).$
Since $\lambda_{1}y(\Omega_{1})=0,~{}y(\Omega)\geq 0,~{}\lambda_{2n-1}\leq\lambda_{2n-3},\lambda_{2n}\leq\lambda_{2n+2}$, we just need to prove that $\lambda_{*}y(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}\lambda_{2n-1}y(\Omega_{2n+1})$ and $\lim\limits_{n\rightarrow+\infty}\lambda_{2n}y(\Omega_{2n})\leq\lambda_{*}y(\Omega_{*}^{-})$.
Lemma 2.24
For $y=Ax^{1}$ and $y=b$, we have
$y(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}y(\Omega_{2n+1}),$
$\lim\limits_{n\rightarrow+\infty}y(\Omega_{2n})\leq y(\Omega_{*}^{-}).$
Proof 2.25
Proof. It is easy to see that $\Omega_{*}=\lim\limits_{n\rightarrow+\infty}\Omega_{2n-1}$ and $\Omega_{*}^{-}=\lim\limits_{n\rightarrow+\infty}\Omega_{2n}$. Now we will prove that $Ax^{1}(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n+1})$ and $\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n})\leq Ax^{1}(\Omega_{*}^{-})$. First we show that
$\lim\limits_{n\rightarrow+\infty}\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{2n+1}}x_{i}^{1}=\sum\limits_{i\in\Omega_{*}}\lim\limits_{n\rightarrow+\infty}A_{i}^{\Omega_{2n+1}}x_{i}^{1}=\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{*}}x_{i}^{1}.$
For any $\varepsilon>0$, $\sum\limits_{i\in\mathcal{M}}|x_{i}^{1}|<+\infty$ shows that $\sum\limits_{i=1}^{+\infty}|x_{2i}^{1}|<+\infty$, and $\exists$ constant $B$ s.t. $|x_{i}^{1}|\leq B$ holds for every $i\in\mathcal{M}$. Then there $\exists I$, s.t. $\sum\limits_{i=I+1}^{+\infty}|x_{2n}^{1}|<\frac{(1-\beta)\varepsilon}{4(1+\beta)}$. Since $\lim\limits_{n\rightarrow+\infty}A_{i}^{\Omega_{2n-1}}=A_{i}^{\Omega_{*}}$, for any $1\leq i\leq I$, there $\exists N(i)$ s.t. $|A_{2i}^{\Omega_{2n-1}}-A_{2i}^{\Omega_{*}}|<\frac{\varepsilon}{2(I+1)B}$ when $n>N(i)$. Similarly, $\exists N(*)$ s.t. $|A_{*}^{\Omega_{2n-1}}-A_{*}^{\Omega_{*}}|<\frac{\varepsilon}{2(I+1)B}$ when $n>N(*)$. Let $N=\max\limits_{i\in\{*,1,\cdots,I\}}N(i)$, then when $n>N$, we have
$|\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{2n-1}}x_{i}^{1}-\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{*}}x_{i}^{1}|$
$\leq$ $\sum\limits_{i\in\Omega_{*}}|A_{i}^{\Omega_{2n-1}}-A_{i}^{\Omega_{*}}||x_{i}^{1}|$
$=$ $|A_{*}^{\Omega_{2n-1}}-A_{*}^{\Omega_{*}}||x_{*}^{1}|+\sum\limits_{i=1}^{I}|A_{2i}^{\Omega_{2n-1}}-A_{2i}^{\Omega_{*}}||x_{2i}^{1}|+\sum\limits_{i=I+1}^{+\infty}|A_{2i}^{\Omega_{2n-1}}-A_{2i}^{\Omega_{*}}||x_{2i}^{1}|$
$\leq$ $\frac{\varepsilon}{2}+\frac{2(1+\beta)}{1-\beta}\cdot\frac{1-\beta}{4(1+\beta)}\cdot\varepsilon$
$=$ $\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.$
So we obtain $\lim\limits_{n\rightarrow+\infty}\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{2n-1}}x_{i}^{1}=\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{*}}x_{i}^{1}$. Thus we have
$\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n-1})=\lim\limits_{n\rightarrow+\infty}\sum\limits_{i\in\Omega_{2n-1}}A_{i}^{\Omega_{2n-1}}x_{i}^{1}$
$\geq$ $\lim\limits_{n\rightarrow+\infty}\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{2n-1}}x_{i}^{1}=\sum\limits_{i\in\Omega_{*}}A_{i}^{\Omega_{*}}x_{i}^{1}=Ax(\Omega_{*}).$
A similar process yields $\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n})\leq Ax^{1}(\Omega_{*}^{-})$. Next we prove
$b(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n-1}),$
$\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n})\leq b(\Omega_{*}^{-}).$
In fact, by Theorem 2.4, if the initial state is $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, $b(\{0\}\cup\Omega)=\frac{1}{1-\beta}-T_{j}^{\Omega^{c}}$. As for $b(\Omega)=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$, when we always active the auxiliary arm, then $\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}=0$. Combining this fact with $b(\Omega)\geq 0$, we have $b(\Omega)=0$. Next, we only need to consider whether state 00 is in $\Omega_{*}$ or not:
(i) If $0\in\Omega_{*}$, then $0\in\Omega_{2n-1}$ holds for every $n$. In this case,
$b(\Omega_{*})=\frac{1}{1-\beta}-T_{j}^{\Omega_{*}^{c}},b(\Omega_{2n-1})=\frac{1}{1-\beta}-T_{j}^{\Omega_{2n-1}^{c}}.$
Since $\lim\limits_{n\rightarrow+\infty}T_{j}^{\Omega_{2n-1}^{c}}=T_{j}^{\Omega_{*}^{c}}$, we have $b(\Omega_{*})=\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n-1})$.
[…]
Based on the previous discussion and Lemma 2.24, we arrive at:
Lemma 2.26
$-\lambda^{*}\in P_{Y^{*}}$
In summary, when the problem satisfies PCL-indexability and Assumption 2.6 holds, we have the following main theorem:
[…]
If the countable-state RMAB problem considered above is PCL-indexable and Assumption 2.6 holds, then the optimal policy for the two-arm problem is the priority policy given by $\{\lambda_{\boldsymbol{\omega_{i}}}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$, i.e., for any $\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, if $\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}$, then the policy gives higher priority to state $\boldsymbol{\omega_{i}}$ than to state $\boldsymbol{\omega_{j}}$. Furthermore, the problem is Whittle indexable, and $\lambda_{\boldsymbol{\omega_{i}}}$ is exactly the Whittle index of the belief state $\boldsymbol{\omega_{i}}$.
Proof 2.28
[…]
3 The Application of PCL in Approximate-State-Space Problems
In this section, we approximate the system state space based on the concept of $T$-step state space $\Omega(T|\boldsymbol{\omega})$ defined in Section 2.3 for effective applications of our PCL framework.
3.1 $T$-Step Approximate State Space
If the repeated states are also taken into account, then $|\Omega(T|\boldsymbol{\omega})|=\frac{(L+1)^{T+1}-1}{L}$. In this case, the number of states will increase exponentially with the number of time steps $T$. However, in practical situations, the phenomenon of state repetitions often occurs due to the long time horizon of the underlying stochastic processes in consideration. In other word, along the process of belief state updates, the updated state may have already appeared within a small number of steps from the start or have a short Euclidean distance to a previous state. In such cases, we can assume that the states encountered within a finite number of steps approximate the true belief state space reasonably well. In this sense, the cardinality of the $T$-step state space can be significantly reduced. Formally, we recursively define the $T$-step approximate state space as follows.
Definition 3.1 ($T$-step approximate state space)
For any $\varepsilon>0$, belief state $\boldsymbol{\omega}$ and belief state space $\Omega$, we say $d(\boldsymbol{\omega},\Omega)\geq\varepsilon$, if $\forall\boldsymbol{\omega}^{\prime}\in\Omega$, we have $||\boldsymbol{\omega}-\boldsymbol{\omega}^{\prime}||\geq\varepsilon$. Define
$\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})=\begin{cases}\{\mathcal{B}_{l_{1}}\cdots\mathcal{B}_{l_{T}}\boldsymbol{\omega}:0\leq l_{1},\cdots,l_{T}\leq L\text{ and}\\ d(\mathcal{B}_{l_{1}}\cdots\mathcal{B}_{l_{T}}\boldsymbol{\omega},\tilde{\Omega}(T-1,\varepsilon|\boldsymbol{\omega}))\geq\varepsilon\}\cup\tilde{\Omega}(T-1,\varepsilon|\boldsymbol{\omega}),&T\geq 2\\ \{\mathcal{B}_{l}\boldsymbol{\omega}:0\leq l\leq L\text{ and }d(\mathcal{B}_{l}\boldsymbol{\omega},\{\boldsymbol{\omega}\})\geq\varepsilon\}\cup\{\boldsymbol{\omega}\},&T=1\end{cases}.$
We call $\tilde{\Omega}(T|\boldsymbol{\omega},\epsilon)$ the $T$-step approximate state space under the initial belief state $\boldsymbol{\omega}$.
Under the definition of $T$-step approximate state space, numerical calculations show that $|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|$ is likely to be much smaller than $|\Omega(T|\boldsymbol{\omega})|$, which will greatly reduce the actual number of states that need to be considered for the problem. The following example illustrates the above concept.
Example 3.2
Set $\textbf{P}=\begin{pmatrix}0.8&0.2\\ 0.2&0.8\end{pmatrix},\textbf{E}=\begin{pmatrix}0.8&0.2\\ 0.2&0.8\end{pmatrix},\textbf{R}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$ and $\boldsymbol{\omega}_{0}=(0.6,0.4)$. Assume the upper limit of norm error $\varepsilon$ is $10^{-3}$. After belief updates based on observations, then the number of belief states $|\tilde{\Omega}(6,10^{-3}|\boldsymbol{\omega})|$ in the 6-step approximate space space and the number of states $|\Omega(6|\boldsymbol{\omega})|$ in the 6-step state space are 309 and 1093. If the case under consideration is $\varepsilon=1e-4$ and $T=10$, we have $|\tilde{\Omega}(10,10^{-4}|\boldsymbol{\omega})|=3550,|\Omega(10|\boldsymbol{\omega})|=88573$. These two examples show that in the approximate sense, the number of states to be considered under the update rule is far less than the number of states in the precise sense, which practically allows us to transform the infinite-state space problem into a finite-state space problem. Furthermore, if the obtained reward is the only observation and the upper limit of norm error $\varepsilon$ is $10^{-3}$, we have $|\tilde{\Omega}(6,10^{-3}|\boldsymbol{\omega})|=97$ and $|\tilde{\Omega}(10,10^{-3}|\boldsymbol{\omega})|=158$.
Extensive numerical simulations give the following two very helpful observations:
Observation 1
For any given $\varepsilon>0$, $\exists T$, such that $\forall t>T$ and $\boldsymbol{\omega}_{t}\in\Omega(t|\boldsymbol{\omega})$, $\exists\boldsymbol{\omega}_{T}\in\tilde{\Omega}(T|\boldsymbol{\omega})$, such that $||\boldsymbol{\omega}_{t}-\boldsymbol{\omega}_{T}||<\varepsilon$, where |||||$\cdot$||| | ⋅ | | is the Euclidean norm.
Observation 2
$\forall\varepsilon>0$, $\lim\limits_{T\rightarrow+\infty}|\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega})|<+\infty$.
Under Observations 1 and 2, we know that if the number of time slots $T$ is sufficiently large, then the new belief states produced after $T$-step updates can be approximated by some belief states in $\tilde{\Omega}(T|\boldsymbol{\omega})$ which is a finite set.
3.2 AG Algorithm for Approximate State Space
Now we are ready to incorporate the approximate state space $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$ into the AG algorithms introduced in Niño-Mora (2002) based on Theorem 2.27 and Observations 1 and 2. Specifically, for the state update, if the resulting belief state $\boldsymbol{\omega}$ does not fall within $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$, then we search for the belief state $\boldsymbol{\omega^{\prime}}$ in $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega_{0}})$ that has the closest Euclidean distance to $\boldsymbol{\omega}$, and use the Whittle index $\lambda_{\boldsymbol{\omega^{\prime}}}$ of $\boldsymbol{\omega^{\prime}}$ as the Whittle index of $\boldsymbol{\omega}$ or adopt a linear interpolation since $\varepsilon$ is small. The detailed algorithm for computing the approximate Whittle index for our RMAB with general observation models is given below.
  Algorithm 1 Approximate Whittle Index Algorithm
1:$\beta\in(0,1),T\geq 1,N\geq 2,1\leq K<N,L,\varepsilon$
2:Initial belief state $\omega^{(n)}(0),\textbf{P}^{(n)},\textbf{E}^{(n)},\textbf{R}^{(n)},n=1,...,N$
3:for $n=1,...,N$ do
4:     Let $\Omega^{(n)}(0,\varepsilon|\boldsymbol{\omega}^{(n)}(0))=\{\boldsymbol{\omega}^{(n)}(0)\}$ and $\Omega^{(n)}(-1,\varepsilon|\boldsymbol{\omega}^{(n)}(0))=\emptyset$
5:     for $t=1,...,T$ do
6:         Let $D_{t-1}=\Omega^{(n)}(t-1,\varepsilon|\boldsymbol{\omega}^{(n)}(0))\backslash\Omega^{(n)}(t-2,\varepsilon|\boldsymbol{\omega}^{(n)}(0))$
7:         Let $\Omega^{(n)}(t,\varepsilon|\boldsymbol{\omega}^{(n)}(0))=\Omega^{(n)}(t-1,\varepsilon|\boldsymbol{\omega}^{(n)}(0))$
8:         for $\omega\in D_{t-1}$ do
9:              for $l=1,...,L$ do Compute $\boldsymbol{\omega^{\prime}}=\mathcal{B}_{l}\boldsymbol{\omega}$
10:                  for $\boldsymbol{\omega^{\prime\prime}}\in\Omega^{(n)}(t,\varepsilon|\boldsymbol{\omega}^{(n)}(0))$ do
11:                       if $||\omega^{\prime}-\omega^{\prime\prime}||>\varepsilon$ then
12:                           Append $\omega^{\prime}$ to $\Omega^{(n)}(t,\varepsilon|\boldsymbol{\omega}^{(n)}(0))$
13:                           break
14:     Compute $P_{(n)}^{1},P_{(n)}^{0},\{R_{\boldsymbol{\omega}_{i}}|\boldsymbol{\omega}_{i}\in\Omega_{1}^{(n)}\}$
15:     Let $\Omega_{1}^{(n)}=\Omega^{(n)}(T,\varepsilon|\boldsymbol{\omega}^{(n)}(0)),k=1,$ compute the sequence $\{A_{\boldsymbol{\omega}_{i}}^{\Omega_{1}^{(n)}}|\boldsymbol{\omega}_{i}\in\Omega_{1}^{(n)}\}$
16:     Compute $\gamma_{\boldsymbol{\omega}_{i}}^{\Omega_{1}^{(n)}}=\frac{R_{\boldsymbol{\omega}_{i}}}{A_{\boldsymbol{\omega}_{i}}^{\Omega_{1}^{(n)}}},\boldsymbol{\omega}_{i}\in\Omega_{1}^{(n)};$
17:     Let $\pi_{1}=\arg\max\{\gamma_{\boldsymbol{\omega}_{i}}^{\Omega_{1}^{(n)}}|\boldsymbol{\omega}_{i}\in\Omega_{1}^{(n)}\},\gamma_{\boldsymbol{\omega}_{\pi_{1}}}^{(n)}=\gamma_{\boldsymbol{\omega}_{\pi_{1}}}^{\Omega_{1}^{(n)}},k=1$
18:     while $|\Omega_{k}^{(n)}|>1$ do
19:         $k=k+1,\Omega_{k}^{(n)}=\Omega_{k-1}^{(n)}\backslash\{\boldsymbol{\omega}_{\pi_{k-1}}\}$
20:         Compute $\gamma_{\boldsymbol{\omega}_{i}}^{\Omega_{k}^{(n)}}=\gamma_{\boldsymbol{\omega}_{i}}^{\Omega_{k-1}^{(n)}}+\left(\frac{A_{\boldsymbol{\omega}_{i}}^{\Omega_{k-1}^{(n)}}}{A_{\boldsymbol{\omega}_{i}}^{\Omega_{k}^{(n)}}}-1\right)(\gamma_{\boldsymbol{\omega}_{i}}^{\Omega_{k-1}^{(n)}}-\gamma_{\boldsymbol{\omega}_{\pi_{k}}}^{\Omega_{k-1}^{(n)}}),\boldsymbol{\omega}_{i}\in\Omega_{k}^{(n)}$
21:         Let $\pi_{k}=\arg\max\{\gamma_{\boldsymbol{\omega}_{i}}^{\Omega_{k}^{(n)}}|\boldsymbol{\omega}_{i}\in\Omega_{k}^{(n)}\},\gamma_{\boldsymbol{\omega}_{\pi_{k}}}^{(n)}=\gamma_{\boldsymbol{\omega}_{\pi_{k}}}^{\Omega_{k}^{(n)}}$
22:     $K=k$
23:     if $\gamma_{\boldsymbol{\omega}_{\pi_{1}}}^{(n)}\geq\cdots\geq\gamma_{\boldsymbol{\omega}_{\pi_{K}}}^{(n)}$ then
24:         $FAIL^{(n)}:=0$
25:     else
26:         $FAIL^{(n)}:=1$
27:if For any $01\leq n\leq N,FAIL^{(n)}=$ 0 then
28:     $FAIL:=0$
29:else
30:     $FAIL:=1$
31:return $\Omega^{(n)}(T,\varepsilon|\boldsymbol{\omega}^{(n)}(0)),\{\gamma_{\boldsymbol{\omega}}^{(n)}|\boldsymbol{\omega}\in\Omega_{1}^{(n)}\},FAIL$, where $1\leq n\leq N$
Remark. In the adaptive-greedy part of this algorithm, we need to input the parameter $\boldsymbol{A}$ of every arm. For the approximate state space, we can solve for all the parameters $A_{\boldsymbol{\omega}}^{\Omega}$ through the dynamic programming equations of $T_{\boldsymbol{\omega}}^{\Omega}$. The output FAIL value is used to judge whether the problem is PCL-indexable w.r.t. $\tilde{\Omega}$. If FAIL=0, the problem is PCL-indexable, otherwise not.
3.3 Numerical Results
In this subsection, we demonstrate the strong performance of our algorithm by numerical simulations. For the calculation of approxiamte state space, we set the maximum number of iterations to 6, and the error limit $\varepsilon$ to $31e-$ 3. For each example, we set the number of arms to 6, and the number of arms activated each time to 1, and other parameters to meet the PCL-indexability. We mainly consider the unit-time-return performance of PCL/Whittle index policy and myopic policy when the RMAB system runs for 200 time slots, using 10000 Monte-Carlo simulation experiments. Figures 2-6 show that our PCL/Whittle index policy achieves a superior performance (specific parameter settings are shown in Table 1-3).
Figure 1: Example 1
Figure 2: Example 2
Figure 3: Example 3
Figure 4: Example 4
Figure 5: Example 5
Figure 6: Example 6
Table 1: Transition Matrices Setting 1
Arm System 1 2 3
1 $\textbf{P}^{(1)}=\begin{pmatrix}0.524&0.221&0.255\\ 0.271&0.277&0.452\\ 0.337&0.281&0.382\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.427&0.212&0.361\\ 0.305&0.320&0.375\\ 0.089&0.383&0.528\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.305&0.333&0.362\\ 0.206&0.331&0.463\\ 0.236&0.179&0.585\end{pmatrix}$
2 $\textbf{P}^{(2)}=\begin{pmatrix}0.671&0.022&0.307\\ 0.396&0.215&0.389\\ 0.095&0.339&0.566\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.504&0.172&0.324\\ 0.355&0.315&0.330\\ 0.129&0.303&0.568\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.477&0.187&0.336\\ 0.391&0.266&0.343\\ 0.009&0.404&0.587\end{pmatrix}$
3 $\textbf{P}^{(3)}=\begin{pmatrix}0.371&0.237&0.392\\ 0.235&0.491&0.274\\ 0.443&0.154&0.403\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.185&0.259&0.556\\ 0.160&0.595&0.245\\ 0.096&0.378&0.526\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.255&0.356&0.389\\ 0.232&0.436&0.332\\ 0.175&0.083&0.742\end{pmatrix}$
4 $\textbf{P}^{(4)}=\begin{pmatrix}0.438&0.298&0.264\\ 0.413&0.460&0.127\\ 0.368&0.233&0.399\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.575&0.279&0.146\\ 0.384&0.353&0.263\\ 0.068&0.201&0.731\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.670&0.081&0.249\\ 0.712&0.102&0.186\\ 0.256&0.531&0.213\end{pmatrix}$
5 $\textbf{P}^{(5)}=\begin{pmatrix}0.610&0.229&0.161\\ 0.183&0.445&0.372\\ 0.565&0.162&0.273\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.261&0.308&0.431\\ 0.215&0.709&0.076\\ 0.422&0.278&0.300\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.394&0.085&0.521\\ 0.261&0.648&0.091\\ 0.467&0.133&0.400\end{pmatrix}$
6 $\textbf{P}^{(6)}=\begin{pmatrix}0.346&0.500&0.154\\ 0.392&0.414&0.194\\ 0.497&0.231&0.272\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.301&0.436&0.263\\ 0.379&0.464&0.157\\ 0.372&0.152&0.476\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.538&0.134&0.328\\ 0.700&0.117&0.183\\ 0.380&0.297&0.323\end{pmatrix}$
Table 2: Transition Matrices Setting 2
Arm System 4 5 6
1 $\textbf{P}^{(1)}=\begin{pmatrix}0.410&0.350&0.240\\ 0.412&0.403&0.185\\ 0.473&0.198&0.329\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.586&0.261&0.153\\ 0.164&0.468&0.368\\ 0.018&0.395&0.587\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.420&0.214&0.366\\ 0.317&0.338&0.345\\ 0.117&0.429&0.454\end{pmatrix}$
2 $\textbf{P}^{(2)}=\begin{pmatrix}0.524&0.221&0.255\\ 0.271&0.277&0.452\\ 0.337&0.281&0.382\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.468&0.461&0.071\\ 0.457&0.448&0.095\\ 0.038&0.550&0.412\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.285&0.362&0.353\\ 0.289&0.373&0.338\\ 0.037&0.378&0.585\end{pmatrix}$
3 $\textbf{P}^{(3)}=\begin{pmatrix}0.364&0.299&0.337\\ 0.346&0.480&0.174\\ 0.322&0.420&0.258\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.263&0.392&0.345\\ 0.317&0.624&0.059\\ 0.247&0.565&0.188\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.538&0.172&0.290\\ 0.379&0.468&0.153\\ 0.408&0.439&0.153\end{pmatrix}$
4 $\textbf{P}^{(4)}=\begin{pmatrix}0.383&0.601&0.016\\ 0.158&0.262&0.580\\ 0.537&0.393&0.070\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.179&0.284&0.537\\ 0.322&0.674&0.004\\ 0.372&0.186&0.442\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.286&0.581&0.133\\ 0.332&0.491&0.177\\ 0.406&0.352&0.242\end{pmatrix}$
5 $\textbf{P}^{(5)}=\begin{pmatrix}0.370&0.154&0.476\\ 0.440&0.178&0.382\\ 0.413&0.231&0.356\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.419&0.253&0.328\\ 0.391&0.267&0.342\\ 0.202&0.393&0.405\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.320&0.306&0.374\\ 0.341&0.381&0.278\\ 0.206&0.568&0.226\end{pmatrix}$
6 $\textbf{P}^{(6)}=\begin{pmatrix}0.262&0.432&0.306\\ 0.294&0.508&0.198\\ 0.250&0.339&0.411\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.387&0.016&0.597\\ 0.270&0.283&0.447\\ 0.136&0.480&0.384\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.163&0.102&0.735\\ 0.053&0.529&0.418\\ 0.121&0.435&0.444\end{pmatrix}$
Table 3: Error and Reward Matrix Setting
Error Matrix Reward Matrix
$\textbf{E}=\begin{pmatrix}0.90&0.05&0.05\\ 0.10&0.80&0.10\\ 0.10&0.05&0.85\end{pmatrix}$ $\textbf{R}=\begin{pmatrix}0&0&0\\ 0&1&0\\ 0&0&2\end{pmatrix}$
Note: the same error and reward matrices were set for each arm in all experiments.
4 Conclusion
In this paper, we generalize the PCL framework for RMAB with finite state spaces to countable state spaces and propose an efficient algorithm to test the PCL-indexability and compute the Whittle index function for the partially observable RMAB with general observation models. As the number of belief states increases rapidly with time, the classic dynamic programming analysis for the RMAB problem suffers from the curse of dimensionality. Based on the achievable region and conservation law analysis studied in the literature for various RMAB problems, this paper is the first to extend the previous theoretical results in this direction to infinite-state scenarios by taking advantage of the theory of infinite-dimensional linear programming. Under a loose regularity assumption, we prove that the PCL-indexability is a sufficient condition for the Whittle indexability. Last, by introducing the concept of approximate state space, we build an algorithm to calculate the approximate Whittle index of each state using a finite-state adaptive greedy algorithm. Numerical results verified the strong performance of the Whittle index policy in our model.
Although we have quite a few theoretical analysis and confirmed the effectiveness of the Whittle index algorithm, there are still some problems to be solved in the future. An important one is that the concept of approximate state space is formed on the basis of a large number of numerical simulations, and we have not yet proved Observations 1 and 2. Another problem is that we only analyzed the optimality of the Whittle index if the problem satisfies the PCL-indexability with respect to some set family $\tilde{\Omega}$, but we have not yet explored other possible sets that may yield PCL to work as well. Furthermore, the strict calculation of Whittle indices for the entire belief state space is still open (even though we only need their order to be correct). All in all, more rooms for exploration of problems with infinite state spaces still await to be unlocked.